Red Royal Entertainment is considering a project that would last for 2 years. The project would involve an initial investment of 171,000 dollars for new equipment that would be sold for an expected price of 143,000 dollars at the end of the project in 2 years. The equipment would be depreciated to 19,000 dollars over 8 years using straight-line depreciation. In years 1 and 2, relevant annual revenue for the project is expected to be 155,000 dollars per year and relevant annual costs for the project are expected to be 37,000 dollars per year. The tax rate is 50 percent and the cost of capital for the project is 7.53 percent. What is the net present value of the project?
Answer #1

Cost of new equipment = $171,000

Equipment Life = 8 years

Salvage Value at the end of 8 year = $19,000

Annual Depreciation on Equipment = (Cost of new equipment- Salvage Value at the end of 8 year

)/ Equipment Life

= (171,000-19,000)/8

= $19,000

.

Book Value of the Equipment at the end of project = 171,000 – 19,000*2

= $133,000

.

Profit on sale of Equipment at the end of project = Sales Price – Book Value

= 143,000 - 133,000

= $10,000

After Tax Salvage Value of Sausage system at the end of project =Sales Price - Profit on sale of Equipment at the end of project*Tax rate)

= 143,000 - 10.000*0.5

= 143,000 - 5,000

=138,000

.

Calculation of NPV of the project:

Year Inititial Investment (1) Annual Revenue (2) Annual Cost (3) Deprecaition(4) Net Profit Before Tax = (5) = (2)-(3)-(4) Tax @ 34% (6) = (5)*0.50 Net Profit after Tax (7) = (5)-(6) Cash Flow from Operations (8) = (7)+(4) After Tax Salavge Value (9) Net Cash Flow (10) = (1)+(8)+(9) DF Working Discounting Factor @7.53% (12) Present Value (13) = (11)*(12)
0                              (171,000)                                     -                                                                            -                                                                            -                                                                                   -                                                      -                                                                                  -                                                                                  -                                                       -                                                                    (171,000)                      1                                                       1                                     (171,000.00)
1                                            -                           155,000                                                                37,000                                                                19,000                                                                        99,000                                           49,500                                                                      49,500                                                                      68,500                                                     -                                                                        68,500 1/1.0753^1 0.929973031                                          63,703.15
2                                            -                           155,000                                                                37,000                                                                19,000                                                                        99,000                                           49,500                                                                      49,500                                                                      68,500                                         138,000                                                                    206,500 1/1.0753^2 0.864849838                                        178,591.49
NPV:                                          71,294.64

.

NPV of the project is $71,294.64.Since , NPV is positive , the project is viable.